I have problems with math, but I have a question for my homework.. how can I do this problem:
Write an equation in slope-intercept:
(-1,5) and (3-3)
.. I hope you answer quickly.
thanks!

Answer by queenofit(120)

Since you do not have the slope, you must first find the slope by using the point-slope formula. Since you have two points use this form. The first set of points are numbered 1 and the second set of points are numbered 2 for this.
that is what I mean when i put the 1 or 2 beside the variable in this equation.
OK??? like this
Y2 - Y1
------- = m=slope your equation would look like this -3-5
X2 - X1 ----- = -8/4=-2
3-(-1)
so the slope of your line is -2
Then use the Point-slope form of the quation of a line
y-y1 = m(x-x1) Your equation would look like y-5=-2(x-(-1)
distribute all the numbers and put in y=mx+b form (slope intercept form)
y-5=-2x-2
y=-2x+5-2
y=-2x+3 slope intercept form.